OpenStudy (anonymous):

A shirt originally cost $40.44, but it is on sale for $34.37. What is the percentage decrease of the price of the shirt? If necessary, round to the nearest percent.

jimthompson5910 (jim_thompson5910):

what's the change in price

OpenStudy (anonymous):

im not sure

jimthompson5910 (jim_thompson5910):

subtract the sale price from the original price

jimthompson5910 (jim_thompson5910):

40.44 - 34.37 = ???

jimthompson5910 (jim_thompson5910):

what do you get?

OpenStudy (anonymous):

6.07?

jimthompson5910 (jim_thompson5910):

yep, so it decreased $6.07 divide this by the original price 6.07/40.44 = ???

OpenStudy (anonymous):

15.01?

jimthompson5910 (jim_thompson5910):

no, how did you get that

jimthompson5910 (jim_thompson5910):

it should be 6.07/40.44 = 0.1500989

jimthompson5910 (jim_thompson5910):

Then you multiply that by 100 to convert it to a percent 0.1500989*100 = 15.00989 So it's roughly 15%

OpenStudy (anonymous):

thankyou

jimthompson5910 (jim_thompson5910):

So what this means is that if you start with $40.44 and you take 15% off, then you'll be left with $34.37